# Anomalous enhancement of Néel order in the $S=1/2$ square lattice Heisenberg model under fictitious magnetic field

### v1

@@ IV Summary
We have studied impacts of the fictitious magnetic flux $\phi$ in the $S=1/2$ Heisenberg model on the square lattice. Within the spin-wave approximation neglecting the magnon interaction, we found that the magnon excitation gap is $\Delta E\sim\sqrt{\phi}$ in agreement with the field theory description in the previous study. As a result, the Néel antiferromagnetic order is enhanced by the application of a fictitious magnetic flux and it shows a scaling behavior $\Delta m_{s}\sim\sqrt{\phi}$. This is reasonable because of the magnon gap, but in contrast to a naive expectation based on the Dzyaloshinskii-Moriya interaction appearing in the lowest-order expansion of the Aharonov-Casher phase. We also performed the DMRG calculations to fully take the magnon interactions and quantum fluctuations into account. Interestingly, we identified a much stronger enhancement of the Néel order with the scaling behavior $\Delta m_{\rm corr}^{z}\sim\phi^{0.136}$ or possibly a fluctuation-induced discontinuous enhancement. The anomalous enhancement compared with the spin-wave approximation clearly suggests importance of the quantum fluctuations of the Heisenberg model which is critical. An alternative picture for the anomalous enhancement was discussed based on the magnetic catalysis for critical systems by rewriting the Heisenberg model as an interacting hard-core boson model. The magnetic catalysis in the present system is stronger than that in any fermion systems and is beyond the standard magnetic catalysis. The quantum spin systems are an interesting platform for the interplay of the Landau quantization and interactions.
@@ Acknowledgements.
The DMRG calculations are performed with the use of ITensor.jl [27]. We thank Shunsuke C. Furuya and Takahiro Misawa for fruitful discussions. This study is supported by JSPS KAKENHI Grant No.22K03513.
@@ Appendix A Details of the spin-wave approximation
$\mathcal{H}_{\rm sw}=JS\sum_{\expectationvalue{i,j}}\left(a^{\dagger}_{i}a_{i}+b^{\dagger}_{i}b_{i}+e^{i\theta_{ij}}a^{\dagger}_{i}b^{\dagger}_{j}+e^{-i\theta_{ij}}a_{i}b_{j}\right),$ (11)
where we have neglected interactions between the bosons. The Hamiltonian in the real space basis can be straightforwardly diagonalized. Since the spin-wave Hamiltonian has the magnetic translation symmetry on a torus, physical quantities are also translationally symmetric. It turns out that translational symmetry is kept in the bulk even for cylinder geometry. Based on this observation, we perform the Fourier transformation in the $y$-direction in the cylinder geometry. The Hamiltonian is written as
@@ Appendix B Finite-size scaling within spin-wave approximation
where $J_{z}/J_{xy}\geqq 1$. The Heisenberg point $J_{z}=J_{xy}$ is SU(2) symmetric and is located at a bicritical point [54], whereas the ground state is Néel ordered along the $z$-axis under the condition $J_{z}>J_{xy}$. Near the critical point $J_{z}=J_{xy}$, we suppose that the deference $\Delta m_{s}$ is described by a scaling function $f$,
@@
where $g=\left(J_{z}-J_{xy}\right)/J_{xy}$ and $\beta$, $\nu$ are the critical exponents at $\phi=0$ for the deference $\Delta m_{s}(g,l_{B}\to\infty)\sim g^{\beta}$ and the correlation length $\xi(g,l_{B}\to\infty)=g^{-\nu}$. Note that we use $l_{B}$ instead of $L$ for the finite-size scaling analysis, since $l_{B}$ is considered to be the dominant length scale rather than the system size $L$ when $l_{B}<L$. In the previous works at $\phi=0$, the spin-wave approximation gives the enhancement of the Néel oder $\Delta m_{s}(J_{z})\propto\sqrt{J_{z}-J_{xy}}$ and energy gap $\Delta E\propto\sqrt{J_{z}-J_{xy}}$ in the leading order [47, 48]. The same critical behaviors have been obtained also by the extensive numerical calculations [44, 45]. These studies give the critical exponents $\beta=0.5$ and also $\nu=0.5$ under the assumption that the dynamical critical exponent is $z=1$ since the correlation length $\xi$ scales as $\xi^{z}\sim\frac{1}{\Delta E}$. Figure 9 shows the finite-$l_{B}$ scaling with the fixed critical exponents $\beta=\nu=1/2$. We see that the numerical data successfully collapse onto a single curve, and the finite-$l_{B}$ scaling indeed holds. Thus the two length scales $l_{B}$ and $L$ are interchangeable within the spin-wave approximation, which may be reasonable. However, this equivalence is broken in the DMRG results where the full quantum fluctuations are taken into account as discussed in the main text.
@@ Appendix C Robustness of power-law fitting for $m^{0}_{\rm corr}$
In the main text, the zero field value of the Néel magnetization $m_{\rm corr}^{0}=0.2985$ is used for the discussions of $\Delta m_{\rm corr}^{z}$. To check robustness of our discussion, we perform the same power-law fitting using a different zero-field value, $m^{0}_{\rm corr}=0.30743$ obtained in the previous quantum Monte Carlo study [41]. Figure 10 shows the logarithmic scale plot of $\Delta m^{z}_{\rm corr}$ versus $\phi$ using $m_{\rm corr}^{0}=0.30743$, similar to Fig. 7 in the main text. The power-law behavior can be seen for $\phi>\phi_{0}$, where the crossover scale is $\phi_{0}=0.05$. The calculation results are summarized in Table 2. The exponent $n$ and the prefactor $C$ are extrapolated to $n\simeq 0.141,C\simeq 0.275$, and they are close to those obtained in the main text. These results indicate that the slight deviation of the background $m_{\rm corr}^{0}$ is negligible in the power-law fitting.

### v2

@@ IV Summary
We have studied impacts of the fictitious magnetic flux $\phi$ in the $S=\frac{1}{2}$ Heisenberg model on the square lattice. Within the spin-wave approximation neglecting the magnon interaction, we found that the magnon excitation gap is $\Delta E\sim\sqrt{\phi}$, which can be qualitatively understood based on the dimensional analysis. As a result, the Néel antiferromagnetic order is enhanced by the application of a fictitious magnetic flux and it shows scaling behavior $\Delta m_{s}\sim\sqrt{\phi}$. This is reasonable because of the magnon gap, but in contrast with a naive expectation based on the Dzyaloshinskii-Moriya interaction appearing in the lowest-order expansion of the Aharonov-Casher phase. We also performed the DMRG calculations to fully consider the magnon interactions and quantum fluctuations. Interestingly, we identified a much stronger enhancement of the Néel order with the scaling behavior $\Delta m_{\rm corr}^{z}\sim\phi^{0.136}$ or possibly a fluctuation-induced discontinuous enhancement. The anomalous enhancement compared with the spin-wave approximation clearly suggests the importance of the quantum fluctuations of the Heisenberg model which is critical. An alternative picture for the anomalous enhancement was discussed based on the magnetic catalysis for critical systems by rewriting the Heisenberg model as an interacting hard-core boson model. The magnetic catalysis in the present system is stronger than that in any fermion systems and is beyond the standard magnetic catalysis. The quantum spin systems are an interesting platform for the interplay of the Landau quantization and interactions.
@@ Acknowledgements.
The DMRG calculations were performed with the use of ITensor.jl [26]. We thank Shunsuke C. Furuya and Takahiro Misawa for fruitful discussions. This study is supported by JSPS KAKENHI Grant No.22K03513.
@@ Appendix A Details of the spin-wave approximation
$\mathcal{H}_{\rm sw}=JS\sum_{\expectationvalue{i,j}}\left([a^{\dagger}_{i}a_{i}+b^{\dagger}_{i}b_{i}+e^{i\theta_{ij}}a^{\dagger}_{i}b^{\dagger}_{j}+e^{-i\theta_{ij}}a_{i}b_{j}\right],$ (11)
where we have neglected interactions between the bosons. The Hamiltonian in the real space basis can be straightforwardly diagonalized. Since the spin-wave Hamiltonian has magnetic translation symmetry on a torus, physical quantities are also translationally symmetric. It turns out that translational symmetry is kept in the bulk even for cylinder geometry. Based on this observation, we perform Fourier transformation in the $y$-direction in the cylinder geometry. The Hamiltonian is written as
@@ Appendix B Finite-size scaling within spin-wave approximation
where $J_{z}/J_{xy}\geqq 1$. The Heisenberg point $J_{z}=J_{xy}$ is SU(2) symmetric and is located at a bicritical point [56], whereas the ground state is Néel ordered along the $z$-axis under the condition $J_{z}>J_{xy}$. Near the critical point $J_{z}=J_{xy}$, we suppose that the difference $\Delta m_{s}$ is described by a scaling function $f$,
@@
where $g=\left(J_{z}-J_{xy}\right)/J_{xy}$ and $\beta$, $\nu$ are the critical exponents at $\phi=0$ for the difference $\Delta m_{s}(g,l_{B}\to\infty)\sim g^{\beta}$ and the correlation length $\xi(g,l_{B}\to\infty)=g^{-\nu}$. Note that we use $l_{B}$ instead of $L$ for the finite-size scaling analysis since $l_{B}$ is considered the dominant length scale rather than the system size $L$ when $l_{B}<L$. In previous works at $\phi=0$, the spin-wave approximation gives the enhancement of the Néel order $\Delta m_{s}(J_{z})\propto\sqrt{J_{z}-J_{xy}}$ and energy gap $\Delta E\propto\sqrt{J_{z}-J_{xy}}$ in the leading order [48, 49]. The same critical behaviors have also been obtained by the extensive numerical calculations [45, 46]. These studies give the critical exponents $\beta=0.5$ and $\nu=0.5$ under the assumption that the dynamical critical exponent is $z=1$ since the correlation length $\xi$ scales as $\xi^{z}\sim\frac{1}{\Delta E}$. Figure 9 shows the finite-$l_{B}$ scaling with the fixed critical exponents $\beta=\nu=\frac{1}{2}$. We see that the numerical data successfully collapse onto a single curve, and the finite-$l_{B}$ scaling indeed holds. Thus the two length scales $l_{B}$ and $L$ are interchangeable within the spin-wave approximation, which may be reasonable. However, this equivalence is broken in the DMRG results where the full quantum fluctuations are considered as discussed in the main text.
@@ Appendix C Robustness of power-law fitting for $m^{0}_{\rm corr}$
In the main text, the zero-field value of the Néel magnetization $m_{\rm corr}^{0}=0.2985$ is used for the discussions of $\Delta m_{\rm corr}^{z}$. To check the robustness of our discussion, we perform the same power-law fitting using a different zero-field value $m^{0}_{\rm corr}=0.30743$ obtained in the previous quantum Monte Carlo study [42]. Figure 10 shows the logarithmic scale plot of $\Delta m^{z}_{\rm corr}$ vs $\phi$ using $m_{\rm corr}^{0}=0.30743$, similar to Fig. 7. The power-law behavior can be seen for $\phi>\phi_{0}$, where the crossover scale is $\phi_{0}=0.05$. The calculation results are summarized in Table 2. The exponent $n$ and the prefactor $C$ are extrapolated to $n\simeq 0.141,C\simeq 0.275$, and they are close to those obtained in the main text. These results indicate that the slight deviation of the background $m_{\rm corr}^{0}$ is negligible in the power-law fitting.